OpenStudy (anonymous):

Anyone help me with this question and ill give the first person to help a medal

OpenStudy (anonymous):

What Question?

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

How can this multiplication problem be written using only improper fractions? five times two and one-eighth equals what value? five over one times sixteen over eight equals what value? five over one times eight over sixteen equals what value? five over one times seventeen over eight equals what value? five over one times eighteen over seven equals what value? --------------------------------------------------------------------------------

OpenStudy (texaschic101):

where is the multiplication problem ?

OpenStudy (anonymous):

right under the question

OpenStudy (texaschic101):

5 x 2 1/8 = 5/1 x 17/8

OpenStudy (texaschic101):

to turn a mixed number into an improper fraction, take the whole number and multiply it by the denominator and then add that to the numerator and put that over the original denominator. watch.... 2 1/8 = (2 x 8 + 1)/8 = 17/8

OpenStudy (texaschic101):

and of course 5 would just be 5/1

OpenStudy (texaschic101):

any questions ?